OpenStudy (anonymous):

solve the equation -7(v+2)+4v+7=8v+9

OpenStudy (lgbasallote):

distribute -7 into v+2 first. what do you get?

OpenStudy (anonymous):

what did you get as an answer because somehow i am getting 11over 1

OpenStudy (aripotta):

you got that after distributing?

OpenStudy (aripotta):

or did you do the whole thing?

OpenStudy (anonymous):

idk what am doing to tell you the truth

OpenStudy (anonymous):

\[-7(v+2)+4v+7=8v+9\]\[-7(v+2)=(8v-4v) + (9-7)\]\[-7(v+2)=4v+2\]\[-14v-14=4v+2\]\[7v+4v=-2-14\]\[11v=-16\]\[v=(-16/11)\] As simple as that :)